# Response of wavelength-shifting and scintillating-wavelength-shifting fibers to ionizing radiation

### v1

@@ 4.1 Beta response
where II is the number of photoelectrons (p.e.) recorded by the SiPM, and $x$ is the distance between the radiation source and the SiPM. $I_{\text{long}}$, $I_{\text{short}}$, $\lambda_{\text{long}}$, and $\lambda_{\text{short}}$ denote long and short components of the light yields and attenuation lengths. It is common in the literature and various applications that only $\lambda_{\text{long}}$ is used as “the fiber attenuation length”. However, at shorter distances a double-exponent function provides a much better description of the light yield [1, 2].
@@ 4.4 Summary and discussion of measurements
For $\beta$ irradiation, the light yield of EJ-160I and EJ-160II is approximately 5.0 and 6.9 times higher than that of BCF-91A, respectively. For $\gamma$ irradiation, the corresponding factors are 5.6 and $7.$ For the $\alpha$ source, the enhancement is more modest, with EJ-160I and EJ-160II yielding factors of 2.9 and 3.6, respectively. This reduction may be attributed to the stronger quenching of the scintillation light for $\alpha$ particles in the polystyrene-based fibers EJ-160I and EJ-160II [20].

### v2

@@ 4.1 Beta response
where $I$ is the number of photoelectrons (p.e.) recorded by the SiPM, and $x$ is the distance between the radiation source and the SiPM. $I_{\text{long}}$, $I_{\text{short}}$, $\lambda_{\text{long}}$, and $\lambda_{\text{short}}$ denote long and short components of the light yields and attenuation lengths. It is common in the literature and various applications that only $\lambda_{\text{long}}$ is used as “the fiber attenuation length”. However, at shorter distances a double-exponent function provides a much better description of the light yield [1, 2].
@@ 4.4 Summary and discussion of measurements
For $\beta$ irradiation, the light yield of EJ-160I and EJ-160II is approximately 5.0 and 6.9 times higher than that of BCF-91A, respectively. For $\gamma$ irradiation, the corresponding factors are 5.6 and 7.7. For the $\alpha$ source, the enhancement is more modest, with EJ-160I and EJ-160II yielding factors of 2.9 and 3.6, respectively. This reduction may be attributed to the stronger quenching of the scintillation light for $\alpha$ particles in the polystyrene-based fibers EJ-160I and EJ-160II [20].